OpenStudy (anonymous):

Select one of the factors of 5x2 + 7x + 2. I'll medalll

OpenStudy (aum):

5x^2 + 7x + 2 = 5x^2 + 5x + 2x + 2 = 5x(x + 1) + 2(x + 1) = (5x+2)(x+1)